# SPECIAL RELATIVITY

### v9

@@ 5.14 Ether Theory and the Aberration of Light Effect
5.14.1 Introductory Remarks
Now we wish to continue in our analysis a little further. We will look for a different way of calculating the aberration effect. Another solution to the accelerated emitter problem will be discussed going back to ether theory. We accept the ether theory in the original form: there is ether, which rules the speed of light. Because the ether is immovable, it is causes anisotropy in every frames moving relative to the ether. Because the non-inertial frame is accelerated with respect to the fixed stars; consequently we must feel an ”ether wind” when measuring light propagation. Firstly we discuss the effect of the ether wind on the light speed. According to the hypothesis of an ether at rest in the initial inertia frame, the velocity of light judged from an accelerated reference system would be $c+u$ for the beam propagating in the same direction as the accelerated system and $-c+u$ for the beam propagating in the opposite direction, where $u=-v$ is the ether velocity in the accelerated reference system. The aberration effect is an effect of the ether-wind on the speed of light (Fig. 25).
Figure 25: Aberration of light in an accelerated frame. According to the ether theory, the ether wind generates anisotropy in the accelerated ( relative to the ether) frame that is responsible for change of radiation direction.
@@ 5.14.2 Lorentz’s Theory of Corresponding States
In 1895, Lorentz derived a first-order version of the so-called ”theorem of the corresponding states” and used it to give a very general explanation of the null results of the first order optical interference experiments. Lorentz’s theorem connects pairs of electromagnetic field configurations, one in a system at rest in the ether and one in a system in uniform motion through the ether, and say that, if one member of the pair is allowed by the laws of electrodynamics, so is the other. The theorem works as follows. The mathematical description of the configuration in the moving system in therms of ”local time” (instead of ”real time” ) is the same as the description of its corresponding state in the system at rest in terms of the real time. Lorentz showed that two field configurations related to one another in many of their observable properties, in particular, they give rise to identical interference patterns. Thus, it is not possible to detect the orbital velocity using interference (e.g. Michelson-Morley) experiments (2).
@@
We now discuss the application of Lorentz’s pre-relativistic theory to the aberration of light effect in a non-inertial frame of reference. Above we considered the problem of aberration of light in the accelerated systems on the basis of theory of special relativity. We described the effect of aberration by working only up to the first order. In fact, the appearance of relativistic effects in optical phenomena does not depend on a large speed of the radiation sources. Light is always a relativistic object.
@@
We have already discussed the asymmetry between initial inertial frame and accelerated frames in this chapter. Here’s what we found out there: We pointed that metric tensor must be a continuous quantity. This problem in special relativity can be adequately treated only by an approach which uses the absolute time coordinatization. The usual Maxwell’s equations are valid in the initial inertial frame where optical system (e.g. configuration of light source and lens, Fig. 13) is at rest and procedure of slow clock transport is used to assign values of time coordinate. We require that metric of a non inertial frame associated with accelerated (with respect to the initial inertial frame) optical system would pass continuosly into Langevin metric. For this purpose, it is sufficient to relate the coordinates and time of the accelerated observer $(x_{n},y_{n},z_{n},t_{n})$ to the coordinates $(x,y,z,t)$ of the inertial observer by the inverse Galilean boost $x_{n}=x-vt,y_{n}=y,z_{n}=z$, $t_{n}=t$. Hence the wave equation transforms into Eq. 16 where coordinates and time are transformed according to the Galilean transformation. The Langevin metric, together with initial condition, describes optical effects in the accelerated frame. Above in this chapter we considered the aberration of light problem in the accelerated frame. The presented approach uses the Fourier transform method. The essential asymmetry is rooted in the electrodynamics equation Eq. 16. As one of the consequences of the crossed term in the wave equation we found the group velocity that is responsible for the aberration. However, there is another satisfactory way of explaining the effect of aberration of light in the accelerated frame. The explanation consists in using a procedure of clock resynchronization. Using Einstein’s clock synchronization we actually select a Lorentz coordinate system for the optical setup at rest in the accelerated frame. In this synchronization we describe the field configuration using usual Maxwell’s equations. The time $t^{\prime}_{n}$ under the Einstein’s synchronization is obtained by substituting $t^{\prime}_{n}=t_{n}-x_{n}v/c^{2}$ in the first order approximation. Clearly, after the resynchronization, the asymmetry is rooted in the initial condition. The plane radiation wavefront of the emitter rotates in the accelerated frame on the angle $-v/c$ after resynchronization.
@@
Lorentz’s treatment of optical setup in an inertial frame moving through the ether is similar to our treatment of such setup in the accelerated (with respect to the fixed stars) frame. Now we must analyze this similarities. One thing is clear: Our initial inertial frame (i.e. inertial frame without accelerational history with respect to the fixed stars) can be interpreted as what Lorentz thought of as a frame at rest in the ether. In the Lorentz’s theory, and more generally in Physics before the development of special relativity, the connection between an observer at rest who used $(x,y,z,t)$ coordinates and an observer bound to the moving body who used $(x_{n},y_{n},z_{n},t)$ coordinates was given by the Galileo relation: $x_{n}=x-vt,y_{n}=y,z_{n}=z,t_{n}=t$. Galilean boost in Lorentz’s theory is empirically equivalent to the smooth tailoring of the metric tensor. In this case the metric of initial inertial frame would pass into the Langevin metric of the accelerated frame. Lorentz of course would not have thought of his theory in this way, since he did not have the resources of Minkowski space-time to do so.
@@
Because of this Galilean boost, the homogeneous wave equation for field in the moving frame, Eq. 16, has not standard form. The main difference consists in the crossed term, which complicates the solution of the equation. To get around of this difficulty, Lorentz observed that simplification is always possible. In solving wave equation, Lorentz found it convenient to introduce new variable $t^{\prime}_{n}=t_{n}-x_{n}v/c^{2}$. It is called ”local time”. The substitution of this time shift in Lorentz theory is a mathematical device used to change the electrodynamics equations into Maxwell’s form.
@@
At first cite, it seems that this Lorentz’s theorem of corresponding states for first order phenomena provides all the guidance we need to exact treatment of aberration of light in the accelerated frame. One could naively expect that Lorentz’s theory would predict in full generality that observation of aberration of light from the earth-based incoherent source would reveal the earth orbital motion, Fig. 24. At this point, a reasonable question arises: why Lorentz concluded from his theorem of corresponding states that no first order effects of the orbital motion of the earth could be detected?
@@ 5.14.3 Analogy with the Sound Theory
A comparison between the Lorentz ether theory and the sound theory might help here. Above in the Section 4.15 we considered the sound emitter problem in the atmosphere frame. We considered the case of a screen, in the atmosphere frame, moving with velocity $v$ along its surface. According to the theory of sound, there is no deviation of the energy transport for sound transmitted through a hole in the moving screen. A plane wave is falling normally on the screen, and generates a transmitted oblique sound beam, Fig. 7. The group velocity is equal to the phase velocity. The wave equation $\nabla^{2}f-\partial^{2}f/\partial(v_{s}t)^{2}=0$ always valid in the atmosphere frame. This equation is not applicable from the viewpoint of an observer at rest with respect to the accelerated frame. Accelerations with respect to the atmosphere frame have an effect on the propagation of sound. In particular, the wave equation transforms into
@@
where coordinates and time are transformed according to the Galilean transformation $x_{n}=x-vt$, $t_{n}=t$. The wave equation needs to be integrated with initial condition for the sound wavefront. Acceleration has no effect on the wavefront orientation. So that wavefront of the emitted sound beam is perpendicular to the $z_{n}$ axis. Because the air is immovable, it is causes anisotropy in every frames moving relative to the atmosphere frame. Consequently we must feel an ”air wind” when measuring sound propagation. According to the wave equation Eq. 18, the velocity of sound judged from an accelerated reference system would be $v_{s}-v$ for the beam propagating in the same direction as the accelerated system and $-v_{s}-v$ for the beam propagating in the opposite direction, where $-v$ is the wind velocity in the accelerated reference system. The aberration effect is an effect of the air wind on the speed of sound (Fig. 25).
@@
Let us now analyze the aberration of sound radiated by a single ”plane-wave” emitter in the atmosphere frame. The wavefront radiated by the plane-wave emitter at given instant time is constructed as an envelope of all the spherical wavelets from the point sources on the emitter that have radiated until this instant. It is not difficult to see that when the emitter is moving at a constant velocity $v$ along the $x$ axis, the wavefront will propagate at the speed $v_{s}$ at the angle $\theta=v/v_{s}$ from the vertical. In fact, due to the motion of the emitter, constructive interference occurs when the projection of phase velocity on the $x$ axis, is equal to the velocity of the emitter. According to classical kinematics, the extra phase chirp $d\phi/dx=k_{x}=vk/v_{s}=v\omega_{s}/v_{s}^{2}$ is introduced. As a consequence of this, the plane wavefront rotates after the emitter acceleration. Then, the radiated sound beam is propagated at the angle $v/v_{s}$ with respect to the $z$axis, yielding the phenomenon of the aberration of light in the atmosphere frame. Now let us return to observations of an accelerated observer. The wave equation needs to be integrated with initial condition for the sound wavefront. Galilean boost has no effect on the wavefront orientation. So that wavefront of the emitted sound beam is tiled at the angle $v/v_{s}$ with respect to the $z_{n}$ axis. The wave number vector of the radiated plane wave is fixed by initial condition. Then we have $k_{x}=vk/v_{s}=v\omega_{s}/v_{s}^{2}$. The non-diagonal wave equation Eq. 18 states that radiated sound beam moves along $x_{n}$ axis with group velocity $d\omega/dk_{x}=-v+v_{s}^{2}k_{x}/\omega_{s}=0$ (see Section 5.11 for more detail).
@@
Thus we obtain the same addition theorem for sound beam velocities as for the particle velocities in the classical physics. So that the two (ether and sound) theories were really just variants of the same classical theory. According to Lorentz, the wave equation of the moving light (sound) emitter in the ether (atmosphere) frame is the same as wave equation of the emitter at rest. As a consequence of this, after the boost we can see that screen acceleration in the ether (atmosphere frame) has no effect on the wavefront orientation and the wave equation. Thus Lorentz concluded that one could not determine the orbital motion of the earth by means of aberration of earth-based source measurements. We come to the conclusion that the Lorentz analysis of the aberration of light does not take into account the fundamental difference between velocity of light and velocity of sound.
@@
When Lorentz formulated his pre-relativistic theory he assumed a space-time background with a well defined geometrical structure: Newtonian space-time. At close look at the physics of this subject shows that there is a real difference between the inertial frame without accelerational history with respect to the fixed stars and the Lorentz classical ether at rest. The main difference between the initial inertial frame and the absolute space of Newton is that special relativity taught us to think in terms of unified space-time model. In special relativity the absolute space and the absolute time of Newton are fused into absolute space-time.
5.15 Bibliography and Notes
1. The aberration of light problem in the accelerated systems is solved with discovery of the essential asymmetry between the non-inertial and the inertial observers. This asymmetry is of the same nature as that of the well-known Sagnac effect [28, 29, 30]. For instance Langevin’s 1921 explanation of the Sagnac effect rested upon the assertion that ”any change of velocity, or any acceleration has an absolute meaning.” [31].
2. For discussion of the relationship between the general accepted way of looking at special relativity and the Lorentz’s ether theory we suggest reading the paper [32].
@@ 6.1 The Corpuscular Model of Light and the Stellar Aberration
It is generally believed that the phenomenon of aberration of light could be interpreted, using the corpuscular model of light, as being analogous of the observation of the oblique fall of raindrops by a moving observer. This is a classical kinematics method to the computation of the stellar aberration used in astronomy for about three hundred years (1). What had to be added in the 20th century, was that the dynamical laws of Newton for light were found to be all wrong, and electromagnetic wave theory had to be introduced to correct them.
As well-known, according to textbooks the physical basis of the stellar aberration is the fact that the velocity of light is finite and changes its direction when seen from another reference frame. It is a consequence of the formula for addition of velocities applied to a light beam when the observer is changing its reference frame.
For an observer on the earth, it is, with respect to the solar referential frame, of about $v=$ 30 km/s, corresponding to the earth motion around the sun. It is sufficient to describe the effect of stellar aberration by working only up to the first order $v/c=10^{-4}$. According to the conventional approach, the study of stellar aberration is intimately connected with the old (Newtonian) kinematics: the Galilean vectorial law of addition of velocities is actually used.
@@ 6.2 Wave Theory of the Stellar Aberration
Most authors treat light propagating through the telescope barrel as rain of photons, and not as a plane wave (2) . Questioning the validity of standard reasoning we argue that a satisfactory treatment of stellar aberration should be based on the coherent wave optics. It can be easily demonstrated that the light produced by a distant star is approximately coherent over a circular area whose diameter, in all practical cases, is much larger than the telescope diameter. Hence, we sample such a tiny portion of the coherent area of the starlight with our telescopes that the waveforms are effectively (flat) plane waves (3).
@@
Let us verify that this assertion is correct. The spatial coherence of a light beam generally has to do with the coherence between two points in the field illuminated by the light source. The meaning of spatial coherence can best to understood with the help of Young’s two-pinhole experiment (see Fig.26). In its elementary sense, the degree of coherence between the two points simply describes the contrast of the interference fringes that are obtained when the two points are taken as secondary sources. Let a source $S$ illuminates the two pinholes $S_{1}$ and $S_{2}$, as shown in Fig.26. The source is perfectly non-coherent. That is to say, no interference fringes can be obtained by placing two pinholes in the plane of the source. It was shown, however, that if the two pinholes are placed far enough away from the non-coherent source, interference fringes of good contrast can be obtained. It is sometimes said that the spatial coherence in light beams increases with distance ”by mere propagation”. It would be nice to find an explanation which is elementary in the sense that we can see what is happened physically.
@@
Suppose that a quasi-monochromatic wave is an incident on an aperture in an opaque screen, as illustrated in Fig. 26. In general, this wave may be partially coherent. The detailed structure of an optical wave undergoes changes as the wave propagates through space. In a similar fashion, the detailed structure of the spatial coherence undergoes changes, and in this sense, the transverse coherence function is said to propagate. Knowing the spatial coherence on the aperture, we wish to find the spatial coherence on the observing screen at distance $z$ beyond the aperture. The stellar radiation is a stochastic object and for any starlight beam there exist some characteristic linear dimension, $\Delta r$, which determines the scale of spatially random fluctuations. Fig.27 illustrates the type of spiky pattern on an aperture in an opaque screen. When $\Delta r\ll d$, the radiation beyond the aperture is partially coherent. This case is shown in Fig.27. Here $\Delta r$ may be estimated as the typical linear dimension of spikes.
[…]
@@
Figure 28: Reciprocal width relation of Fourier transform pair. Free space basically act as a Fourier transformation. This means that the field in the far zone is, phase factor and proportionality factor aside, the spatial Fourier transform of the field at the source plane.
A star is an incoherent source with the scale of spatially random fluctuations $\lambdabar\Delta r\sim\lambdabar$. Such a source emits radiation in all directions. Also, an angular spectrum of the starlight contains spikes with the typical width $(kd)^{-1}$, where $d$ is the star diameter. In the case of astronomical observations by telescope, each spike on the earth’s surface has a finite thickness which is the same order as the spatial coherence length of the starlight. It can be said that the coherent area of the observed starlight would be of the order of $z/(kd)$, where $z$ is the distance between the star and earth. The star Bradley chose was Draconis, which is one of the nearest stars. In this case the diameter of the coherent area on the earth surface can be estimated to be about 100 m.

### v10

@@ 5.13 Ether Theory and the Aberration of Light Effect
Now we wish to continue in our analysis a little further. Another solution to the accelerated emitter problem will be discussed going back to ether theory. We accept the ether theory in its original form: there is ether, which rules the speed of light. Because the ether is immovable, it causes anisotropy in every frame moving relative to the ether. Because the non-inertial frame is accelerated with respect to the fixed stars; consequently we must feel an ”ether wind” when measuring light propagation. Firstly we discuss the effect of the ether wind on the light speed. According to the hypothesis of an ether, the velocity of light judged from an accelerated reference system would be $c+u$ for the beam propagating in the same direction as the accelerated system and $-c+u$ for the beam propagating in the opposite direction, where $u=-v$ is the ether velocity in the accelerated reference system. The aberration effect is an effect of the ether-wind on the speed of light (Fig. 23).
Figure 23: Aberration of light in an accelerated frame. According to the ether theory, the ether wind generates anisotropy in the accelerated ( relative to the ether) frame that is responsible for the change of radiation direction.
@@ 5.13.1 Lorentz’s Theory of Corresponding States
In 1895, Lorentz derived a first-order version of the so-called ”theorem of the corresponding states” and used it to give a very general explanation of the null results of the first-order optical interference experiments. Lorentz’s theorem connects pairs of electromagnetic field configurations, one in a system at rest in the ether and one in a system in uniform motion through the ether, and says that, if one member of the pair is allowed by the laws of electrodynamics, so is the other. The theorem works as follows. The mathematical description of the configuration in the moving system in terms of ”local time” (instead of ”real time” ) is the same as the description of its corresponding state in the system at rest in terms of real time. Lorentz showed that two field configurations related to one another in many of their observable properties, in particular, they give rise to identical interference patterns. Thus, it is not possible to detect the orbital velocity using interference (e.g. Michelson-Morley) experiments  [^27] For discussion of the relationship between the generally accepted way of looking at special relativity and Lorentz’s ether theory we suggest reading the paper [32]..
@@
We now discuss the application of Lorentz’s pre-relativistic theory to the aberration of light effect in a non-inertial frame of reference. Above we considered the problem of aberration of light in the accelerated systems based on the theory of special relativity. We described the effect of aberration by working only up to the first order. We already know that light is always a relativistic object.
@@
We have already discussed the asymmetry between initial inertial frame and accelerated frames in this chapter. Here’s what we found out there: We pointed out that the metric tensor must be a continuous quantity. This problem in special relativity can be adequately treated only by an approach that uses absolute time coordinatization. The usual Maxwell’s equations are valid in the initial inertial frame where the optical system (e.g. configuration of light source and lens, Fig. 13) is at rest and the procedure of slow clock transport is used to assign values of time coordinate. We require that the metric of a noninertial frame associated with an accelerated (with respect to the initial inertial frame) optical system pass continuously into the Langevin metric. For this purpose, it is sufficient to relate the coordinates and time of the accelerated observer $(x_{n},y_{n},z_{n},t_{n})$ to the coordinates $(x,y,z,t)$ of the inertial observer by the inverse Galilean boost $x_{n}=x-vt,y_{n}=y,z_{n}=z$, $t_{n}=t$. Hence the wave equation transforms into Eq. 17 where coordinates and time are transformed according to the Galilean transformation. The Langevin metric, together with the initial condition, describes optical effects in the accelerated frame. Above in this section, we considered the aberration of light problem in the accelerated frame. The presented approach uses the Fourier transform method. The essential asymmetry is rooted in the electrodynamics equation Eq. 17. As one of the consequences of the crossed term in the wave equation, we found the group velocity that is responsible for the aberration. However, there is another satisfactory way of explaining the effect of aberration of light in the accelerated frame. The explanation consists of using a procedure of clock resynchronization. Using Einstein’s clock synchronization, we select a Lorentz coordinate system for the optical setup at rest in the accelerated frame. In this synchronization, we describe the field configuration using Maxwell’s equations. The time $t^{\prime}_{n}$ under Einstein’s synchronization is obtained by substituting $t^{\prime}_{n}=t_{n}-x_{n}v/c^{2}$ in the first-order approximation. Clearly, after the resynchronization, the asymmetry is rooted in the initial condition. The plane radiation wavefront of the emitter rotates in the accelerated frame on the angle $-v/c$ after resynchronization.
@@
At first sight, Lorentz’s treatment of optical setup in an inertial frame moving through the ether is similar to our treatment of such setup in the accelerated (with respect to the fixed stars) frame. Now we must analyze this similarity. One thing is clear: Our initial inertial frame (i.e. inertial frame without accelerational history with respect to the fixed stars) can be interpreted as what Lorentz thought of as a frame at rest in the ether. In Lorentz’s theory, and more generally in Physics before the development of special relativity, the connection between an observer at rest who used $(x,y,z,t)$ coordinates and an observer bound to the moving body who used $(x_{n},y_{n},z_{n},t)$ coordinates was given by the Galileo relation: $x_{n}=x-vt,y_{n}=y,z_{n}=z,t_{n}=t$. Galilean boost in Lorentz’s theory is empirically equivalent to the smooth tailoring of the metric tensor. In this case, the metric of the initial inertial frame would pass into the Langevin metric of the accelerated frame. Lorentz of course would not have thought of his theory in this way, since he did not have the resources of Minkowski space-time to do so.
@@
Because of this Galilean boost, the homogeneous wave equation for the field in the moving frame, Eq. 17, has no standard form. The main difference consists in the crossed term, which complicates the solution of the equation. To get around of this difficulty, Lorentz observed that simplification is always possible. In solving wave equation, Lorentz found it convenient to introduce a new variable $t^{\prime}_{n}=t_{n}-x_{n}v/c^{2}$. It is called ”local time”. The substitution of this time shift in Lorentz theory is a mathematical device used to change the electrodynamics equations into Maxwell’s form.
@@
At first sight, it seems that Lorentz’s theorem of corresponding states for first-order phenomena provides all the guidance we need to exact treatment of the aberration of light in the accelerated frame. One could naively expect that Lorentz’s theory would predict in full generality that observation of aberration of light from the earth-based incoherent source would reveal the earth’s orbital motion, Fig. 22. At this point, a reasonable question arises: why did Lorentz conclude from his theorem of corresponding states that no first-order effects of the orbital motion of the earth could be detected?
@@ 5.13.2 Analogy with the Sound Theory
A comparison between the Lorentz ether theory and the sound theory might help here. Above in Section 4.13, we considered the sound emitter problem in the atmosphere frame. We considered the case of a screen, in the atmosphere frame, moving with velocity $v$ along its surface. According to the theory of sound, there is no deviation of the energy transport for sound transmitted through a hole in the moving screen. A plane wave falls normally on the screen and generates a transmitted oblique sound beam, Fig. 7. The group velocity is equal to the phase velocity. The diagonal wave equation $\nabla^{2}f-\partial^{2}f/\partial(v_{s}t)^{2}=0$ always valid in the atmosphere frame. This equation is not applicable from the viewpoint of an observer at rest with respect to the accelerated frame. Accelerations with respect to the atmosphere frame affect the propagation of sound. In particular, the wave equation transforms into
@@
where coordinates and time are transformed according to the Galilean transformation $x_{n}=x-vt$, $t_{n}=t$. The wave equation needs to be integrated with the initial condition for the sound wavefront. Acceleration does not affect the wavefront orientation. So wavefront of the emitted sound beam is perpendicular to the $z_{n}$ axis. Because the air is immovable, it causes anisotropy in every frame moving relative to the atmosphere frame. Consequently, we must feel an ”air wind” when measuring sound propagation. According to the wave equation Eq. 19, the velocity of sound judged from an accelerated reference system would be $v_{s}-v$ for the beam propagating in the same direction as the accelerated system and $-v_{s}-v$ for the beam propagating in the opposite direction, where $-v$ is the wind velocity in the accelerated reference system. The aberration effect is an effect of the air wind on the speed of sound (Fig. 23).
@@
Let us now analyze the aberration of sound radiated by a single ”plane-wave” emitter in the atmosphere frame. The wavefront radiated by the plane-wave emitter at a given instant time is constructed as an envelope of all the spherical wavelets from the point sources on the emitter that have radiated until this instant. It is not difficult to see that when the emitter is moving at a constant velocity $v$ along the $x$ axis, the wavefront will propagate at the speed $v_{s}$ at the angle $\theta=v/v_{s}$ from the vertical. In fact, due to the motion of the emitter, constructive interference occurs when the projection of phase velocity on the $x$ axis, is equal to the velocity of the emitter. According to classical kinematics, the extra phase chirp $d\phi/dx=k_{x}=vk/v_{s}=v\omega_{s}/v_{s}^{2}$ is introduced. As a consequence of this, the plane wavefront rotates after the emitter acceleration. Then, the radiated sound beam is propagated at the angle $v/v_{s}$ with respect to the $z$axis, yielding the phenomenon of the aberration of light in the atmosphere frame. Now let us return to observations of an accelerated observer. The wave equation needs to be integrated with the initial condition for the sound wavefront. Galilean boost does not affect the wavefront orientation. So the wavefront of the emitted sound beam is tiled at the angle $v/v_{s}$ with respect to the $z_{n}$ axis. The wave number vector of the radiated plane wave is fixed by the initial condition. Then we have $k_{x}=vk/v_{s}=v\omega_{s}/v_{s}^{2}$. The non-diagonal wave equation Eq. 19 states that the radiated sound beam moves along $x_{n}$ axis with group velocity $d\omega/dk_{x}=-v+v_{s}^{2}k_{x}/\omega_{s}=0$ (see Section 5.10 for more detail).
@@
Thus we obtain the same addition theorem for sound beam velocities as for the particle velocities in classical physics. According to Lorentz, the wave equation of the moving light (sound) emitter in the ether (atmosphere) frame is the same as the wave equation of the emitter at rest. Thus Lorentz concluded that one could not determine the orbital motion of the earth using aberration of earth-based source measurements. So the two (ether and sound) theories were just variants of the same classical theory. We conclude that the Lorentz analysis of the aberration of light does not take into account the fundamental difference between the velocity of light and the velocity of sound.
@@
When Lorentz formulated his pre-relativistic theory he assumed a space-time background with a well-defined geometrical structure: Newtonian space-time. A close look at the physics of this subject shows that there is a real difference between the inertial frame without an accelerational history with respect to the fixed stars and the Lorentz classical ether at rest. The main difference between the initial inertial frame and the absolute space of Newton is that special relativity taught us to think in terms of a unified space-time model. In special relativity, the absolute space and the absolute time of Newton are fused into absolute space-time.
@@ 6.1 The Corpuscular Model of Light and the Stellar Aberration
It is generally believed that the aberration of light could be interpreted, using the corpuscular model of light, as being analogous to observing the oblique fall of raindrops by a moving observer. This is a classical kinematics method to the computation of the stellar aberration used in astronomy for about three hundred years  [^28] The phenomenon of the annual apparent motion of celestial objects about their locations, named stellar aberration, was discovered by Bradley in 1727, who also explained it employing the corpuscular model of light [33].. What had to be added in the 20th century, was that the dynamical laws of Newton for light were found to be all wrong, and electromagnetic wave theory had to be introduced to correct them.
As well-known, according to textbooks the physical basis of the stellar aberration is the fact that the velocity of light is finite and changes its direction when seen from another reference frame. It is a consequence of the formula for the addition of velocities applied to a light beam when the observer is changing its reference frame.
For an observer on the earth, it is, with respect to the solar referential frame, of about $v=$ 30 km/s, corresponding to the earth’s motion around the sun. It is sufficient to describe the effect of stellar aberration by working only up to the first order $v/c=10^{-4}$. According to the conventional approach, the study of stellar aberration is intimately connected with the old (Newtonian) kinematics: the incorrect assumption that an earth-based observer and sun-based observer have a common 3-space is used.
@@ 6.2 Wave Theory of the Stellar Aberration
Most authors treat light propagating through the telescope barrel as rain of photons, and not as a wave  [^29] There is a widespread belief that the rays of light coming from the star are falling on the telescope tube and not interact with its sides. To quote French [4]: ”Regarding light as being composed of rain of photons, we can easily calculate the change in apparent direction of a distant object such a star.” Similar statements can also be found in recently published textbooks. To quote Rafelski [5]: ”We consider a light ray originating in a distant star. … The following discussion is addressing the observation of well-focusing light rays and not (spherical) plane wave light. … This picture of the experimental situation is accurate since the light emitted by a star consists of an incoherent flux of photons produced in independent atomic processes.” This conclusion is incorrect. We emphasize that the design of a telescope optical system is based on the classical theory of diffraction (originated directly from the classical electromagnetic theory). The resolution of a telescope is limited by the diffraction of light waves and the telescope always influences the measured radiation (due to unavoidable diffraction of a starlight by the telescope aperture). In a well-corrected optical system and a circular pupil, the size of the Airy disc of diffraction pattern (i.e. the size of the image of a point source) is inversely proportional to the pupil’s diameter. . Questioning the validity of standard reasoning we argue that a satisfactory treatment of stellar aberration should be based on coherent wave optics. It can be easily demonstrated that the light produced by a distant star is approximately coherent over a circular area whose diameter, in all practical cases, is much larger than the telescope diameter. Hence, we sample such a tiny portion of the coherent area of the starlight with our telescopes that the waveforms are effectively (flat) plane waves  [^30] A star is a completely incoherent source. The character of the mutual intensity function produced by an incoherent source is fully described by the Van Cittert-Zernike theorem [34]. Any star can be considered as very far away from the sun. In all cases of practical interest, telescopes are situated in the far zone of the (distant star) source. Consider a star like Sirius, which is one of the nearest stars. The coherent area of light observed from Sirius has a diameter of about 6 m. This correlation was observed by Brown and Twiss in 1956 [35]..
@@
Let us verify that this assertion is correct. The spatial coherence of a light beam generally has to do with the coherence between two points in the field illuminated by the light source. The meaning of spatial coherence can best to understood with the help of Young’s two-pinhole experiment (see Fig.24). In its elementary sense, the degree of coherence between the two points simply describes the contrast of the interference fringes that are obtained when the two points are taken as secondary sources. Let a source $S$ illuminate the two pinholes $S_{1}$ and $S_{2}$, as shown in Fig.24. The source is perfectly non-coherent. That is to say, no interference fringes can be obtained by placing two pinholes in the plane of the source. It was shown, however, that if the two pinholes are placed far enough away from the non-coherent source, interference fringes of good contrast can be obtained. It is sometimes said that the spatial coherence in light beams increases with distance ”by mere propagation”. It would be nice to find an explanation that is elementary in the sense that we can see what is happening physically.
@@
Suppose that a quasi-monochromatic wave is an incident on an aperture in an opaque screen, as illustrated in Fig. 24. In general, this wave may be partially coherent. The detailed structure of an optical wave changes as the wave propagates through space. In a similar fashion, the detailed structure of the spatial coherence changes, and in this sense, the transverse coherence function is said to propagate. Knowing the spatial coherence on the aperture, we wish to find the spatial coherence on the observing screen at a distance $z$ beyond the aperture. The stellar radiation is a stochastic object and for any starlight beam, there exists some characteristic linear dimension, $\Delta r$, which determines the scale of spatially random fluctuations. Fig.25 illustrates the type of spiky pattern on an aperture in an opaque screen. When $\Delta r\ll d$, the radiation beyond the aperture is partially coherent. This case is shown in Fig.25. Here $\Delta r$ may be estimated as the typical linear dimension of spikes.
@@
Figure 26: Reciprocal width relation of Fourier transform pair. Free space basically acts as a Fourier transformation. This means that the field in the far zone is, phase factor and proportionality factor aside, the spatial Fourier transform of the field at the source plane.
[…]
A star is an incoherent source with the scale of spatially random fluctuations $\lambdabar\Delta r\sim\lambdabar$. Such a source emits radiation in all directions. Also, an angular spectrum of the starlight contains spikes with the typical width $(kd)^{-1}$, where $d$ is the star diameter. In the case of astronomical observations by telescope, each spike on the earth’s surface has a finite thickness which is the same order as the spatial coherence length of the starlight. It can be said that the coherent area of the observed starlight would be of the order of $z/(kd)$, where $z$ is the distance between the star and the earth. The star Bradley chose was Draconis, one of the nearest stars. In this case, the diameter of the coherent area on the earth’s surface can be estimated to be about 100 m.